# Capacity-Aware Inference: Mitigating the Straggler Effect in Mixture of Experts

### v1

@@ 5 Experiments
Table 1: Performance comparison across different capacity factors and selection metrics (i.e., Order, Reverse Order, Random, and Score). The baseline operates without capacity constraints, represented as +$\infty$+ ∞. Due to inherent randomness, we report the average performance over multiple random seeds.
@@
Random 39.6 72.1 57.8 68.3 73.8 45.8 74.2 45.2 59.6
Score 43.2 76.1 56.3 69.9 73.4 52.9 77.1 47.5 62.1
Order 1.0 36.0 60.2 53.4 62.6 69.6 38.7 58.0 36.9 51.9
@@
Random 34.0 63.1 55.2 60.8 70.2 40.5 66.9 35.7 53.3
Score 40.4 71.5 57.0 64.3 71.9 45.4 72.0 39.7 57.8
5.1 Token Drop for High-load Experts
@@
(a) Single MoE Layer
(b) End-to-End Model
Figure 4: Speedup achieved by capacity-aware inference compared to the baseline without capacity constraints.
@@
We next explore the efficiency improvements achieved by imposing expert capacity. Specifically, we employ distributed inference using eight A30 GPUs, utilizing an 8-way Data Parallelism (DP) and 8-way Expert Parallelism (EP) strategy through the Megatron-LM framework Shoeybi et al. (2019). In Mixtral-8$\times$7B-Instruct model, each GPU hosts a single expert, whereas, in models like OLMoE-Instruct, multiple experts must be deployed on a single GPU (e.g., eight experts per GPU) due to GPU resource constraints.
As illustrated in Figure 4, imposing constraints on expert capacity considerably accelerates inference across the four tested MoE models, in comparison to the baseline model without capacity limitations. The enhanced efficiency of each MoE layer (Figure 4 (a)) contributes to faster end-to-end inference (Figure 4 (b)). For instance, when the capacity factor $\gamma$ is set to 1.5, this configuration achieves a 1.94$\times$ speedup for a single MoE module and a 1.37$\times$ speedup for the entire end-to-end model. Furthermore, it is evident that the efficiency improves as the capacity decreases.
Notably, the degree of acceleration is influenced by the numerical relationship between the number of experts and GPUs in Expert Parallelism. In Mixtral-8×8$\times$8 ×7B-Instruct, one-to-one deployment maximizes the effectiveness of capacity-aware inference. Conversely, in models where multiple experts share a single GPU, the acceleration gains are relatively modest. This occurs because the aggregated load from multiple experts diminishes the proportion of reduced load, which is achieved by limiting the straggler expert. Additionally, we believe that further system optimizations or allocating more GPUs for expert distribution would further enhance the potential of capacity-aware inference.
@@
Given that expert capacity enforces MoE layers to discard overflowed tokens, we next establish the relationship between expert capacity and the corresponding number of dropped tokens. Specifically, for a capacity factor $\gamma$, the total proportion of dropped tokens is given by:
@@
Figure 5 visualizes the number of dropped tokens across different capacity factors for various test datasets, with a more detailed illustration provided in Appendix C. Although the most overloaded expert receives much more tokens than the expected number of tokens $\bar{N}$, regulating the maximum capacity has a limited impact on the overall number of accommodated tokens (e.g., reducing it by less than 20% under a capacity factor of 2.0), thereby maintaining competitive performance even after discarding overflow tokens. Moreover, dropping a small proportion of overflowed tokens can significantly reduce the latency caused by overloaded experts (e.g., dropping 12% of overloaded tokens promotes the inference speed by 94% in Mixtral-8$\times$7B-Instruct), highlighting the effectiveness of capacity-constrained inference in improving both performance and efficiency.
Table 2: Comparison of Expert Drop, Token Drop and Token Reroute. The capacity factor $\gamma$ is set to 2.0 for OLMoE and DeepSeek-V2-Lite, and 1.5 for Qwen1.5-MoE-Chat and Mixtral-8×7B-Instruct. In Expert Drop, we only skip one out of eight experts for Mixtral-8$\times$7B-Instruct, and 10% of the lowest load experts for other models.
@@
Token Reroute 47.8 85.0 71.8 83.0 88.6 71.5 87.6 70.2 64.6 74.5
5.2 Token Reroute to Low-load Experts
Besides the experts overloaded with tokens, some low-load experts receive only a few tokens, raising important questions: Are these low-load experts redundant and removable, or should they be leveraged to balance token allocation? Recent works Lu et al. (2024); He et al. (2024) remove less important experts to improve efficiency, while our proposed Token Reroute increases their utilization by redistributing tokens for a more balanced assignment. Next, we investigate the significance of low-load experts and validate the effectiveness of Token Reroute.
@@ 5.2 Token Reroute to Low-load Experts
Effectiveness of Token Reroute
We examine the effectiveness of utilizing low-load experts by rerouting overflowed tokens to them (i.e., Token Reroute) instead of simply discarding these tokens to meet the target capacity. Comparing Token Reroute with Token Drop, redistributing excess tokens to low-load experts enhances performance, yielding a 0.9% improvement in the average performance of Qwen1.5-MoE-Chat. Furthermore, considering the performance degradation observed in Expert Drop, our findings highlight the crucial role of low-load experts in maintaining model effectiveness.
@@
5.3 Ablation Study
Model-Specific Imbalanced Property
We explore the imbalance property in various models, such as OLMoE, DeepSeek-V2-Lite and Qwen1.5-MoE, which differ in both architecture (e.g., depth and width) and training strategies (e.g., training from scratch Muennighoff et al. (2024); DeepSeek-AI et al. (2024a) vs. training after upcycling Jiang et al. (2024); Team (2024b)).
On the one hand, our findings in Appendix B reveal different training strategies result in significantly varying levels of imbalance. Specifically, MoE models trained from scratch exhibit a much higher degree of imbalance. For instance, OLMoE and DeepSeek-V2-Lite experience peak expert-wise token allocations exceeding $5\bar{N}$, whereas Qwen1.5-MoE and Mixtral are upcycled from dense language models, maintain a more balanced distribution, with peak expert-wise allocations staying below $3\bar{N}$. This is because upcycling initializes experts with the same parameters Komatsuzaki et al. (2023), while experts are randomly initialized when training from scratch, leading to a greater imbalance in token assignments Lo et al. (2024).
@@ 5.3 Ablation Study
Beyond the specific capacity values presented in Table 1, we further investigate a wide range of capacity factors in Figure 6, spanning from 0.0 to 3.0. We exclude values exceeding 3.0, as their performance closely aligns with capacity-agnostic scenarios. By analyzing the performance changes when decreasing the capacity factor, we find that setting $\gamma$ to 2.0 is sufficient to maintain performance comparable to the original models. This means allowing each expert to handle utmost twice the average number of tokens maintains the performance of capacity-free inference. However, maintaining the performance becomes challenging under low capacity factors (e.g., 1.0), as high-load experts experience significant token drops.
Figure 7: Ablation study on the number of iterations for Token Reroute.
Rerouting Round
The Token Reroute Algorithm iteratively discourages tokens from selecting overflowed experts while encouraging their rerouting to other available experts, promoting a more balanced token distribution. As iterations progress, lower-score token-to-expert mappings are increasingly considered, further refining the redistribution process. Figure 7 illustrates the results across different iterations. In the first iteration, only excess tokens are dropped, while token rerouting becomes more prominent in subsequent rounds. The first two rounds of token rerouting lead to significant performance improvements, after which performance begins to saturate. We attribute this saturation to the selection of additional low-score token-to-expert mappings, which have a smaller impact compared to the dominant token assignments.
@@ 6 Conclusion
In this paper, we first identify the imbalanced token-to-expert assignment in Mixture of Experts (MoE) and introduce the Straggler Effect in MoE inference, where the highest-load expert becomes the efficiency bottleneck, dictating overall latency. To address this issue, we propose Capacity-Aware Token Drop, which alleviates excessive loads on individual experts, and Capacity-Aware Token Reroute, which enhances the utilization of underutilized experts. Our findings and proposed methods offer valuable insights and effective strategies for improving MoE inference efficiency.
Limitations
Despite the progress we have made, our work still has certain limitations. First, our study focuses on a subset of MoE models, including OLMoE, Qwen1.5-MoE, DeepSeek-V2-Lite, and Mixtral. A promising direction for future work is to extend our analysis and methods to other architectures to further validate their effectiveness across a broader range of MoE models. Second, our proposed Capacity-Aware Token Drop and Capacity-Aware Token Reroute strategies could be integrated into the training process, potentially enhancing both performance and efficiency.
@@ Appendix A Implementation Details
We mainly focus on lightweight MoE models (less than 20B parameter budget). We conduct experiments on OLMoE Muennighoff et al. (2024), Qwen1.5-MoE Team (2024b), DeepSeek-V2-Lite DeepSeek-AI et al. (2024a) and Mixtral Jiang et al. (2024), due to their competitive performance and widespread adoption.
@@
To evaluate model performance, we report normalized zero-shot or few-shot accuracy on the LM-Harness benchmark. The number of shots for each task is detailed in Table 3, which includes multiple tasks: ARC-C Clark et al. (2018), BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021), OBQA Mihaylov et al. (2018), PIQA Bisk et al. (2019), RTE Wang et al. (2019), WinoGrande ai2 (2019) and GSM8K Cobbe et al. (2021). The evaluation code is based on EleutherAI’s LM Harness framework Gao et al. (2023).
@@
Appendix B Layer-wise Expert Load

### v3

@@ 5 Experiments
5.1 Token Drop for High-load Experts
Table 1: Performance comparison across different capacity factors and selection metrics (i.e., Order, Reverse Order, Random, and Score). The baseline operates without capacity constraints, represented as +$\infty$+ ∞. We report the average performance over multiple random seeds.
@@ 5.1 Token Drop for High-load Experts
Random 39.6 72.1 53.8 68.3 73.8 45.8 74.2 45.2 59.1
Score 44.8 77.5 55.2 70.8 74.3 53.4 78.6 50.0 63.1
Order 1.0 36.0 60.2 52.2 62.6 69.6 38.7 58.0 36.9 51.8
@@
Random 34.0 63.1 53.2 60.8 70.2 40.5 66.9 35.7 53.1
Score 41.6 76.0 53.4 69.9 73.2 50.4 77.1 47.8 61.1
@@
Figure 4: Speedup of a single MoE layer compared to the baseline without capacity constraints, achieved through two capacity-aware inference methods: Token Drop and Expanded Drop.
Figure 5: End-to-End Model Speedup
Figure 6: Breakdown Analysis on OLMoE
@@
We next explore the efficiency improvements achieved by imposing expert capacity. Specifically, we employ distributed inference using eight H20 GPUs, utilizing an 8-way Data Parallelism (DP) and 8-way Expert Parallelism (EP) strategy through the Megatron-LM framework [36]. Notably, in Mixtral-8$\times$7B-Instruct model, each GPU hosts a single expert, whereas, in models like OLMoE-Instruct, multiple experts must be deployed on a single GPU (e.g., eight experts per GPU) due to GPU resource constraints.
As illustrated in Figure 4, imposing constraints on expert capacity through Token Drop and Expanded Drop, considerably accelerates inference across the four tested MoE models, in comparison to the baseline model without capacity limitations. The enhanced efficiency of each MoE layer (Figure 4) contributes to faster end-to-end model inference (Figure 6). Moreover, as the capacity factor $\gamma$ decreases, capacity-aware inference methods achieve significantly greater acceleration.
Notably, the efficacy of acceleration is influenced by the numerical relationship between the total experts and the engaged GPUs in Expert Parallelism. As illustrated in Figure 6, for Mixtral-8×8$\times$8 ×7B-Instruct, deploying a single expert per GPU maximizes the effectiveness of capacity-aware inference. In this configuration, Token Drop and Expanded Drop achieve end-to-end model speedups of 1.87 $\times$ and 1.85$\times$, respectively, with $\gamma=1.5$. Conversely, deploying a greater number of experts on a single GPU results in more modest acceleration gains, as evidenced by the “8E/GPU” (OLMoE-Instruct and Deepseek-V2-Lite) the “10E/GPU” (Qwen1.5-MoE-Chat) in Figure 4 and Figure 6. This is because the aggregated load from multiple experts diminishes the proportion of reduced load, which is achieved by limiting the straggler expert. Therefore, it is anticipated that allocating more GPUs for expert distribution, thereby reducing the number of experts per GPU, would enhance the acceleration effect of capacity-aware inference.
The breakdown analysis presented in Figure 6 demonstrates that our proposed capacity-aware inference methods substantially reduce the duration of expert computation, permutation and communication, while preserving a comparable cost for gate processing. Notably, the duration of permutation and communication increases when tokens are expanded across a range of global experts. This is due to the increased communication workload required to transmit expanded global tokens across various GPU devices. Consequently, these results underscore the necessity of restricting the expanded tokens to be processed by local experts.
@@
Given that expert capacity enforces MoE layers to discard overflowed tokens, we next establish the relationship between expert capacity and the corresponding number of dropped tokens. For a capacity factor $\gamma$, the total proportion of dropped tokens is given by:
@@
Figure 7 visualizes the number of dropped tokens across different capacity factors for various test datasets, with a more detailed illustration provided in Appendix F. Although the most overloaded expert receives much more tokens than the expected number of tokens $\bar{N}$, regulating the maximum capacity has a limited impact on the overall number of accommodated tokens, thereby maintaining competitive performance even after discarding overflow tokens. Moreover, dropping a small proportion of overflowed tokens can significantly reduce the latency caused by overloaded experts (e.g., dropping 12% of overloaded tokens promotes the inference speed by 85% in Mixtral-8$\times$7B-Instruct), highlighting the efficacy of capacity-aware inference in improving both performance and efficiency.
Table 2: Comparison of Expert Drop, Token Drop and Expanded Drop. The capacity factor $\gamma$ is set to 2.0 for OLMoE and DeepSeek-V2-Lite, and 1.5 for Qwen1.5-MoE-Chat and Mixtral-8×7B-Instruct. For Expert Drop, each forward pass skips one out of eight experts for Mixtral-8$\times$7B-Instruct, and the bottom 10% of lowest load experts for other models.
@@
Expanded Drop 47.8 85.0 71.8 83.0 88.6 71.5 87.6 70.2 64.6 74.5
5.2 Expanded Drop to Low-load Experts
Besides the experts overloaded with tokens, some low-load experts receive only a few tokens, raising important questions: Are these low-load experts redundant and removable, or should they be leveraged to balance token allocation? Recent works [29, 19] remove less important experts to improve efficiency, while our proposed Expanded Drop increases their utilization by redistributing tokens for a more balanced assignment. Next, we investigate the significance of low-load experts and validate the effectiveness of Expanded Drop.
@@ 5.2 Expanded Drop to Low-load Experts
Figure 8: Normalized expert load after Token Drop and Expanded Drop, with the capacity ratio set to $\gamma=1.5$.
Effectiveness of Expanded Drop
We examine the effectiveness of utilizing low-load experts by Expanded Drop instead of simply discarding these tokens to meet the target capacity. Comparing Expanded Drop with Token Drop, redistributing excess tokens to low-load experts enhances performance, yielding a 0.9% improvement in the average performance of Qwen1.5-MoE-Chat. Furthermore, considering the performance degradation observed in Expert Drop, our findings highlight the crucial role of low-load experts in maintaining model effectiveness.
Expanded Drop overselects experts for each token to expand the selection scope to prompt a more balanced token-expert assignments. As shown in Figure 8, increasing the overselection ratio $m$ allows tokens to consider more candidate experts after being dropped from overflowed ones, thereby improving low-load expert utilization and balancing the expert load.
5.3 Extension to Multi-modal Mixture of Experts
In addition to applying capacity-aware inference to MoE models for language tasks, we also explore its effectiveness in multi-modal MoE settings. Specifically, we evaluate the OLMoE based MolmoE [12], across multi-modal benchmarks, including MME [15], MMBench [28], and SEED-Bench [27].
Table 3: Capacity-aware inference for Multi-modality MoE models under different routing strategies. “Percep.” and “Cognit.” denote Perception and Cognition, respectively. $\gamma$ is set to 1.0.
Method Strategy Percep. Cognit.
Baseline 1358.1 269.6
Token Drop Uniform 1248.4 245.4
Expanded Drop 1307.6 273.6
Token Drop Text First 1114.2 214.4
Expanded Drop 1163.6 241.3
Token Drop Image First 1346.5 288.9
Expanded Drop 1362.1 297.1
Given that the input sequence contains tokens from multiple modalities, we first investigate different token dropping strategies. Specifically, we first treat all tokens equally and drops those with the lowest scores (“Uniform”). Beyond this, considering the redundancy often found in image tokens, we also experiment with a strategy that prioritizes dropping image tokens before selectively removing text tokens (“Image First”). For comparison, we also consider drop text tokens first (“Text First”). As shown in Table 3, on the MME benchmark, this image-first strategy yields improved performance, highlighting the benefit of prioritizing dropping image tokens for the load balance in multi-modal MoE models.
Figure 9: Multi-modal token assignments across different experts.
Figure 10: Comparison on MMBench across six multimodal capabilities in Appendix B.
Given the redundancy of image tokens and their large proportion in multi-modal tasks, we further investigate more aggressive capacity factors for Token Drop and Expanded Drop using the “Image First” strategy. Figure 10 demonstrates the effectiveness of Capacity-Aware Inference under low capacity constraints (i.e., $\gamma=0.5$). This is largely due to the high redundancy in image tokens [4], which allows a higher dropping ratio without significantly affecting performance. Meanwhile, the dominance of image tokens enables the use of very low capacity factors without significantly affecting text token retention, as illustrated in Figure 10. Therefore, dropping image tokens at higher ratios leads to more balanced token assignments and substantially improved inference efficiency.
5.4 Ablation Study
Model-Specific Imbalanced Property
We explore the imbalance property in various models, such as OLMoE, DeepSeek, Qwen and Mixtral, which differ in both architectures (e.g., depth and width) and training strategies (e.g., training from scratch [31, 10] vs. training after upcycling [22, 39]).
On the one hand, our findings in Appendix E reveal different training strategies result in significantly varying levels of imbalance. Specifically, MoE models trained from scratch exhibit a much higher degree of imbalance. For instance, OLMoE and DeepSeek-V2-Lite experience peak expert-wise token allocations exceeding $5\bar{N}$, whereas Qwen1.5-MoE and Mixtral are upcycled from dense language models, maintain a more balanced distribution, with peak expert-wise allocations staying below $3\bar{N}$. This is because upcycling initializes all experts with identical parameters [24], reducing divergence and promoting balanced training in the early stages.
@@ 5.4 Ablation Study
Beyond the specific capacity values presented in Table 1, we further investigate a wide range of capacity factors in Figure 11, spanning from 0.0 to 3.0. We exclude values exceeding 3.0, as their performance closely aligns with capacity-agnostic scenarios. By analyzing the performance changes when decreasing the capacity factor, we find that setting $\gamma$ to 1.5 is sufficient to maintain performance comparable to the original models. However, maintaining performance becomes challenging under excessively low capacity factors, as high-load experts are forced to drop a significant number of tokens.
@@ 6 Conclusion
In this paper, we identify the issue of imbalanced token-to-expert assignment in Mixture of Experts (MoE) models and introduce the Straggler Effect during inference, where high-load experts become efficiency bottlenecks and dictate overall latency. To address this problem, we propose Capacity-Aware Token Drop, which mitigates expert overload by enforcing strict capacity constraints. Additionally, to better utilize underloaded experts, we present Capacity-Aware Expanded Drop, which allows tokens to select additional experts on the same device while still respecting capacity limits, thereby improving expert utilization. Our findings and proposed methods offer valuable insights and effective strategies for improving MoE inference efficiency.
@@ Appendix A Implementation Details
We mainly focus on lightweight MoE models (less than 20B parameter budget). We conduct experiments on OLMoE [31], Qwen1.5-MoE [39], DeepSeek-V2-Lite [10] and Mixtral [22], due to their competitive performance and widespread adoption.
@@
To evaluate model performance, we report normalized zero-shot or few-shot accuracy on the LM-Harness benchmark. The number of shots for each task is detailed in Table 4, which includes multiple tasks: ARC-C [7], BoolQ [6], HellaSwag [43], MMLU [20], OBQA [30], PIQA [2], RTE [40], WinoGrande [1] and GSM8K [8]. The evaluation code is based on EleutherAI’s LM Harness framework [16].
@@
Appendix B Multi-Modal Tasks
In the scope of this paper, multimodal tasks refer to those involving both vision and language modalities. We evaluate model performance using three representative benchmarks: MME, MMBench, and SEED-Bench, each targeting different aspects of multimodal understanding and reasoning.
MME benchmark evaluates vision-language models along two dimensions: perception, which tests visual grounding and recognition, and cognition, which assesses reasoning abilities such as counting and relational understanding. It provides a fine-grained analysis of multimodal understanding.
MMBench is a comprehensive benchmark designed to assess the general multimodal understanding ability of vision-language models. It evaluates model performance across six core capabilities: Coarse Perception (CP), Fine-grained Perception—including single-instance (FP-S) and cross-instance (FP-C), Attribute Reasoning (AR), Logical Reasoning (LR), and Relational Reasoning (RR). By covering both perception and reasoning-oriented tasks, MMBench provides detailed insights into the strengths and limitations of VLMs across diverse multimodal scenarios.
SEED-Bench is a large-scale benchmark for evaluating the generative comprehension of Multimodal Large Language Models (MLLMs) across both image and video modalities. It includes 19K human-annotated multiple-choice questions spanning 12 evaluation dimensions, enabling objective and efficient assessment without human or GPT intervention. SEED-Bench reveals model limitations and maintains a public leaderboard to support fair comparison and future research.
Component Content Token Count
Image [Uncaptioned image] 576
Text Prompt Is this artwork titled virgin and child with sts catherine, cecilia, barbara, and ursula? Please answer yes or no. 31
Total 607
Table 5: An example multi-modal query in MME benchmark, showing the dominant proportion of image tokens compared to text tokens.
As shown in Table 5, these tasks typically introduce a large number of image tokens. When faced with imbalanced token-to-expert assignments, dropping redundant image tokens significantly improves load balancing. Moreover, due to the high redundancy among image tokens, dropping a portion of them has minimal impact on model performance.
As in MME and MMBench, the Image-First variants of Token Drop and Expanded Drop also exhibit consistent effectiveness on SEED-Bench [27], maintaining strong performance even under low capacity factors such as $\gamma=0.5$. In addition to the redundancy in image tokens, Figure 10 shows that text tokens constitute only a small portion of the total token assignments. This allows the regulation of $\gamma$ to retain almost all text tokens under the Image-First strategy.
Method $\gamma$ Inst. Attr. Inst. ID Inst. Interact. Inst. Loc. Inst. Count Scene Spatial Text Reasoning Overall
Baseline $\infty$ 74.2 71.4 58.8 62.8 57.0 73.5 49.6 72.6 76.4 68.7
Token Drop $0.5$ 70.4 67.8 60.8 57.8 51.8 71.5 43.5 58.3 71.3 64.9
Expanded Drop 71.2 67.3 57.7 58.8 53.6 71.0 45.2 64.3 71.9 65.5
Token Drop $1.0$ 73.6 70.2 58.8 62.6 56.9 72.7 48.1 61.9 73.4 68.0
Expanded Drop 73.8 70.7 59.8 63.5 56.7 73.1 49.6 70.2 74.6 68.4
Token Drop $1.5$ 73.5 71.2 58.8 62.9 57.1 73.1 48.7 65.5 74.6 68.3
Expanded Drop 73.7 71.0 61.9 64.7 57.4 73.3 49.3 71.4 76.1 68.7
Table 6: Token Drop and Expanded Drop strategies for multi-modal MoE models evaluated on SEED-Bench. Abbreviations: Inst. Attr. = Instance Attributes; Inst. ID = Instance Identification; Inst. Interact. = Instance Interaction; Inst. Loc. = Instance Localization; Inst. Count = Instance Counting.
Appendix C Pseudocode for Token Drop and Expanded Drop
We present the detailed pseudo-code for the algorithm implementations, as shown in Algorithm 1 and Algorithm 2. Both of these two methods adopt the selection metrics of “Score.”
Algorithm 1 Token Drop
input_tokens, num_tokens, num_experts, $k$, $\gamma$
[…]
scores $\leftarrow\text{Softmax}(\text{logits},\text{dim}=-1)$
topk_scores, topk_indices $\leftarrow\text{TopK}(\text{scores},\text{k}=k,\text{dim}=1)$
[…]
$\triangleright$ Drop Token
expert_capacity $\leftarrow\gamma\cdot\frac{\text{num\_tokens}\times k}{\text{num\_experts}}$
_, capacity_indices $\leftarrow$ TopK(topk_masked_scores, k=expert_capacity, dim=0, sorted=False)
[…]
return final_scores, final_map
Algorithm 2 Expanded Drop
input_tokens, num_tokens, num_experts, $k$, $\gamma$, local_expert_id_list
[…]
scores $\leftarrow\text{Softmax}(\text{logits},\text{dim}=-1)$
topk_scores, topk_indices $\leftarrow\text{TopK}(\text{scores},\text{k}=k,\text{dim}=1)$
$\triangleright$ Expand
local_indices $\leftarrow$ torch.tensor(local_expert_id_list).repeat(num_tokens, 1)
[…]
local_scores $\leftarrow$ scores[:, local_expert_id_list]
[…]
$\triangleright$ Drop Token
expert_capacity $\leftarrow\gamma\cdot\frac{\text{num\_tokens}\times k}{\text{num\_experts}}$
_, capacity_indices $\leftarrow$ TopK(expanded_masked_scores, k=expert_capacity, dim=0, sorted=False)
[…]
return final_scores, final_map
Appendix D Maximum Expert Selection
Expanded Drop employs overselection and dropping to not only ensure load balance but also improve the utilization of underloaded experts. Although this process may allow some tokens to select more than $k$ experts, Table 7 shows that enforcing a strict $k$-expert limit is unnecessary: allowing additional expert selections has the potential of enhancing representational capacity, while rigid constraints introduce redundant computations.
Table 7: Ablation study on limiting the maximum number of $k$ selected experts. “w/max” and “w/o max” indicate runs with and without this constraint, respectively. $\gamma$ is set to 1.0.
$m$ Method OBQA PIQA RTE WinoGrande BoolQ ARC-C HellaSwag MMLU Avg.
$2k$ w/ max 42.4 75.8 53.2 68.6 72.7 50.3 77.1 47.6 61.0
w/o max 42.2 75.9 53.4 69.7 73.1 50.3 77.0 47.8 61.2
$3k$ w/ max 42.0 75.6 53.4 69.6 72.7 50.3 77.0 47.4 61.0
w/o max 42.0 75.6 53.8 69.8 72.9 50.3 77.1 47.6 61.1
Appendix E Layer-wise Expert Load